# latexfontFont shape

### v1

@@ 5.3. ExpTime-hardness of the $\omega$-explorability problem
Before giving the detailed proof of this result, we will give a quick sketch to convey the main ideas.
@@ 5.3.1. Proof sketch of Theorem 31
We give a quick summary of the proof in this section. The full proof can be found in Section 5.3.2. The main idea will be to reduce the acceptance problem of a PSpace alternating Turing machine (ATM) to the $\omega$-explorability problem of some automaton that we build from the machine. This reduction is an adaptation of the one from [BDG+19] showing ExpTime-hardness of the NFA population control problem (defined in Section 4.1).
The computation of an ATM can be seen as a game between two players, who respectively aim for acceptance and rejection of the input. These players influence the output by choosing the transitions when facing a non-deterministic choice, that can belong to either one of them.
Let us first describe the automaton built in [BDG+19]. In that reduction, the choices made by the ATM players are translated into choices for Determiniser and Spoiler. The automaton has two main blocks: one dedicated to keeping track of the machine’s configuration, which we call Config, and another focusing on the simulation of the ATM choices, which we call Choices. In Config, there is no non-determinism: the tokens move following the transitions of the machine given as input to the automaton. In Choices, Determiniser can pick a transition by sending his token to the corresponding state, while Spoiler uses letters to indicate which transition of the ATM he wants to follow.
The automaton constructed this way will basically read a sequence of runs of the ATM. At each run, some tokens must be sent into both blocks. Reaching an accepting state of a run lets Spoiler send some tokens from Choices to his target state $\bot$. He can then restart with the remaining tokens until all are in the target state $\bot$. This process will ensure a win for Spoiler if he has a winning strategy in the ATM game. If he does not, then Determiniser can use a strategy ensuring rejection in the ATM game to avoid the configurations where he loses tokens, provided he starts with enough tokens.
This equivalence between acceptance of the ATM and the automaton being a positive instance of the PCP provides the ExpTime-hardness of their problem.
In our setup, getting rid of tokens one by one is not enough: Spoiler needs to be able to target a specific token and send it to the target state (which is now the rejecting state $\bot$) in one run. If he can do that, repeating the process for every token, without omitting any, ensures his win. If he cannot, then Determiniser has a strategy to pick a specific token and preserving it from $\bot$, and therefore wins.
This is why we adapt our reduction to allow Spoiler to target a specific token, no matter where it chooses to go. To do so, we change the transitions so that winning a run additionally lets Spoiler send every token from Config into $\bot$. With that and the fact that he can already target a token in Choices, we get a winning strategy for Spoiler when the ATM is accepting.
If the ATM is rejecting, Spoiler is still able to send some tokens to $\bot$, but he no longer has that targeting ability, which is how Determiniser is able to build a strategy preserving a specific token to win. To ensure the sustainability of this method, Determiniser needs to keep $\omega$ additional tokens following his designated token, so that he always has $\omega$ tokens to spread into the gadgets every time a new run starts.
Overall, we are able to compute in polynomial time from the ATM a safety automaton that is $\omega$-explorable if and only if the ATM rejects its input. Since acceptance of a polynomial space ATM is known to be ExpTime-hard, we obtain Theorem 31.
@@ 5.3.2. Detailed proof of Theorem 31
We take an alternating Turing machine $\mathcal{M}=(\Sigma_{\mathcal{M}},Q_{\mathcal{M}},\Delta_{\mathcal{M}},q_{0}^{\mathcal{M}},q_{f}^{\mathcal{M}})$ with $Q_{\mathcal{M}}=Q_{\exists}\uplus Q_{\forall}$ and $\Delta_{\mathcal{M}}\subseteq Q_{\mathcal{M}}\times\Sigma_{\mathcal{M}}\times Q_{\mathcal{M}}$. It can be seen as a game between two players: existential ($\exists$) and universal ($\forall$). On a given input, the game creates a run by starting from $q_{0}^{\mathcal{M}}$, and letting $\exists$ (resp. $\forall$) solve the non-determinism in states from $Q_{\exists}$ (resp. $Q_{\forall}$) by picking a transition from $\Delta_{\mathcal{M}}$ compatible with the current letter of the input word. Player $\exists$ wins if the play reaches the accepting state $q_{f}^{\mathcal{M}}$, and $w$ is accepted if and only if $\exists$ has a winning strategy on $w$. We assume that $\mathcal{M}$ uses polynomial space $P(n)$ in the size $n$ of its input, i.e. the winning strategies can avoid configurations with tape longer than $P(n)$. We also fix an input word $w\in(\Sigma_{\mathcal{M}})^{*}$.
We will assume for simplicity that $\Sigma_{\mathcal{M}}=\{0,1\}$ and that the machine alternates between existential and universal states, starting with an existential one (meaning that $q_{0}\in Q_{\exists}$ and the transitions are either $Q_{\exists}\rightarrow Q_{\forall}$ or $Q_{\forall}\rightarrow Q_{\exists}$). In our reduction, this will mean that we give the choice of the transition alternatively to Spoiler (playing $\exists$) and Determiniser ($\forall$).
We create a safety automaton $\mathcal{A}=(Q,\Sigma,q_{0},\Delta,\bot)$ with:
  • $Q=Q_{\mathcal{M}}\uplus\textit{Pos}\uplus\textit{Mem}\uplus\textit{Trans}\uplus\{q_{0},\texttt{store},\bot,\top\}$ where:
    Pos $=[1,P(n)]$ (representing where is the reading head on $\mathcal{M}$’s tape)
    Mem =$=\{m_{b,i}\mid b\in\{0,1\},i\in[1,P(n)]\}$ (representing the current content of $\mathcal{M}$’s tape)
    Trans =$=\{E\}\cup\{A_{t}\mid t\in\Delta_{\mathcal{M}}\}$ (encoding players’ choices)
  • $\begin{array}[t]{ll}\Sigma=&\{a_{t,p}\mid t\in\Delta_{\mathcal{M}}\textrm{ and}p\in[1,P(n)]\}\\ &\uplus\{\texttt{init},\texttt{end},\texttt{restart},\texttt{win}\}\\ &\uplus\{\texttt{check}_{q}\mid q\in Q_{\mathcal{M}}\}\\ &\uplus\{\texttt{check}_{b,i}\mid(b,i)\in\{0,1\}\times[1,P(n)]\}\end{array}.$
  • $\bot$ is a rejecting sink state: a run is accepting if and only if it never reaches this state.
@@
We now define the transitions in $\Delta$. The states $\top$ and $\bot$ are both sinks ($\top$ accepting and $\bot$ rejecting). We then describe all transitions labelled by the letter $a_{t,p}$ with $p\in\textit{Pos}$ and $t=(q,q^{\prime},b,b^{\prime},d)\in\Delta_{\mathcal{M}}$, where $q$ and $q^{\prime}$ are the starting and destination states of $t$, $b$ and $b^{\prime}$ are the letters read and written at the current head position, and $d\in\{L,R\}$ is the direction taken by the head. These transitions are:
  • $q\rightarrow q^{\prime}$.
  • $p\rightarrow p^{\prime}$ with $p^{\prime}=p+1$ if $d=R$, or $p-1$ if $d=L$. It goes to $\top$ if $p^{\prime}\notin[1,P(n)]$.
  • $m_{b,p}\rightarrow m_{b^{\prime},p}$, and $m_{b^{\prime\prime},p^{\prime\prime}}\rightarrow m_{b^{\prime\prime},p^{\prime\prime}}$ for any $b^{\prime\prime}$ and any $p^{\prime\prime}\neq p$.
  • $E\rightarrow A_{t^{\prime}}$ for any transition $t^{\prime}$.
  • $A_{t}\rightarrow E$.
  • $q^{\prime\prime}\rightarrow\top$ for any $q^{\prime\prime}\neq q$.
  • $m_{\neg b,p}\rightarrow\top$ ($\neg b$ is the boolean negation of $b$).
  • $p^{\prime}\rightarrow\top$ for any $p^{\prime}\neq p$.
@@
This is represented in Figure 6.
@@
Figure 6. Transitions for $a_{t,p}$, where $t=(q,q^{\prime},b,b^{\prime},d)$, $p^{\prime}$ is the position at direction $d$ from $p$, and $q^{\prime\prime}$, $p^{\prime\prime}$, and $t^{\prime}$ are different from $q,p,t$ respectively.
The first three bullet points manage the evolution of the configuration of $\mathcal{M}$. The next two deal with the alternation between players, and the next three punish Spoiler if the transition is invalid (the check letters will handle the case where Determiniser is the one giving an invalid transition). The last one saves the tokens that are not chosen for the transition.
The other letters give the following transitions.
  • init goes from $q_{0}$ to the states $E$, $q_{0}^{\mathcal{M}}$, and $1\in\textit{Pos}$, and also to the states $m_{b,p}$ corresponding to the initial content of the tape, i.e. all $m_{b,p}$ such that $b$ is the $p$-th letter of $w$ (or 00 if $i>|w|$).
  • end labels transitions from any non-accepting state of $\mathcal{M}$ to $\top$, from store to $q_{0}$, and from any other state to $\bot$.
  • $\texttt{check}_{q}$ creates a transition from $A_{t}$ to $\bot$ for any $t\in\Delta$ starting from $q$. It also creates a transition from $q$ to $\top$. Any other state is sent back to $q_{0}$. Intuitively, playing that letter means that $q$ is not the current state and that any transition starting from $q$ is invalid.
  • $\texttt{check}_{b,p}$ creates a transition from $A_{t}$ to $\bot$ for any $t\in\Delta$ reading $b$ on the tape. It also creates transitions from any $j\in\textit{Pos}\setminus\{p\}$ and from $m_{b,p}$ to $\top$. Any other state is sent to $q_{0}$. Intuitively, playing that letter means that the current head position is $p$, and that its content is not $b$, so any transition reading $b$ is invalid.
To summarize, the states of $\mathcal{A}$ can be seen as two blocks, apart from $q_{0}$, $\top$ and $\bot$: those dealing with the configuration of $\mathcal{M}$ ($Q_{\mathcal{M}}$, Pos and Mem), and those from the gadget of Figure 7 which deal with the alternation and non-deterministic choices.
@@
We will now prove that $\mathcal{A}$ is $\omega$-explorable if and only if the Turing machine $\mathcal{M}$ rejects the word $w$. Let us first assume that $w\in\mathcal{L}(\mathcal{M})$. There is a winning strategy $\sigma_{\exists}$ for $\exists$ in the alternating Turing machine game, and Spoiler will use that strategy in the explorability game to win against $\omega$ tokens. He will consider that the tokens are labelled by integers, and always target the smallest one that is not already in $\bot$. He proceeds as follows.
@@
  • After such a play, Determiniser can move tokens to any state $A_{t}$. If there is more than one occupied state, Spoiler picks the one containing the current target token (possibly imaginary).
    • If that state corresponds to an invalid transition (wrong starting state or wrong tape content at the current head position), then Spoiler plays the corresponding check letter. Formally, if the target token (not the imaginary one, since Spoiler can avoid invalid transitions for that one) is in $\mathcal{A}_{t}$, Spoiler plays $\texttt{check}_{q}$ if the starting state $q$ of $t$ does not match the current state of the tape (given by Lemma 32), or $\texttt{check}_{b,p}$ if the current head position is $p$ and does not contain $b$. In both cases, the target token is sent to $\bot$ with no other token reaching $\top$ (by Lemma 32). This sends us back to the first step, but with an updated target.
    • If the state instead corresponds to a valid transition, then Spoiler can play the corresponding $a_{t,p}$, where $p$ is the current head position (again, given by Lemma 32), then go back to the previous step (where there are tokens in $E$).
  • If no invalid transition is reached, the run eventually gets to an accepting state of $\mathcal{M}$ because $\sigma_{\exists}$ is winning. This corresponds to a stage where Spoiler can safely play end to get rid of the target token along with all tokens outside of store, by sending them to $\bot$ (the only reason not to play end would be the existence of tokens in non-accepting states of $Q_{\mathcal{M}}$). This sends us back to the first step, but with an updated target. Notice that if there was a virtual target token, we will always reach this event, and send the real target token (located in $Q_{\mathcal{M}}$ or Pos or Mem) in $\bot$.
@@
Conversely, if there is a winning strategy $\sigma_{\forall}$ for the universal player in the alternation game on $\mathcal{M}(w)$, then we can build a winning strategy for Determiniser in the $\omega$-explorability game. This strategy is more straightforward than the previous one, as we can focus on the tokens sent to $E$ (while still populating each state when init is played, but these other tokens follow a deterministic path until the next init).
Determiniser will initially choose a specific token, called leader. He then sends $\omega$ tokens to every reachable state when Spoiler plays init, with the leader going to $E$. Determiniser then moves the tokens in the leader’s state according to $\sigma_{\forall}$. Spoiler cannot send the leader to $\bot$, since the only way to do that would be using the letter end, but this would immediately ensure the win for Determiniser, as there will always be some token in non-accepting states of $\mathcal{M}$ (because $\sigma_{\forall}$ is winning), and those tokens would be sent to $\top$ upon playing end. This means that Spoiler has no way to send the leader to $\bot$ without losing the game, and therefore Determiniser wins.
Note that with that strategy, Spoiler can still safely send some tokens to $\bot$ by playing the wrong transition, which sends the tokens following the leader to store, then some well-chosen check letter to send the remaining ones to $\bot$. However, Determiniser will start the next run with still $\omega$ tokens, including the leader. This is why the choice of a specific leader is important, as it can never be safely sent to $\bot$.

### v2

@@ 5.3. ExpTime-hardness of the $\omega$-explorability problem
Before giving the detailed proof of this result, we briefly sketch its main ideas.
@@ 5.3.1. Proof sketch of Theorem 28
We provide here an intuitive overview of the proof; the full construction is given in Section 5.3.2. The goal is to reduce the acceptance problem of a PSpace alternating Turing machine (ATM) $\mathcal{M}$ to the $\omega$-explorability problem of a safety automaton $\mathcal{A}$. The computation of an ATM can be viewed as a game between two players, $\exists$ and $\forall$, who respectively aim for acceptance and rejection of the input. Each state of the machine belongs to one of the players, who chooses the outgoing transition when in control of that state. The input is accepted by the ATM if $\exists$ has a winning strategy in this game.
The automaton $\mathcal{A}$ will simulate the configuration of $\mathcal{M}$ within its set of reachable states, while reading input letters that describe only local transitions of $\mathcal{M}$. Our construction is an adaptation of the reduction from [BDG+19], which established ExpTime-hardness of the NFA population control problem (defined in Section 4.1). We first recall the principles of that reduction and then explain how we modify them to obtain our result.
Alphabet and input words.
Unlike classical encodings of Turing-machine computations as sequences of configurations, the alphabet here does not contain entire configurations, nor even the current control state of $\mathcal{M}$. Each letter instead specifies a local instruction of the machine:
  • a transition $t=(q,a,q^{\prime},a^{\prime},d)$ of the ATM, and
  • a tape position $i$ indicating on which cell this transition is applied.
Hence, a word over this alphabet represents a possible sequence of transition applications of $\mathcal{M}$, describing at each step where the transition occurs, but leaving the global configuration implicit.
How configurations are represented.
The configuration of $\mathcal{M}$ is encoded by the set of states currently reachable in $\mathcal{A}$. Each state of $\mathcal{A}$ represents either the content of one tape cell, the current position of the head, or the current control state of the machine. Additional states are used for technical purposes, such as recording which player is in control. When a letter $(t,i)$ is read, the automaton updates this distributed encoding: it changes the symbol of cell $i$, moves the head marker, and updates the relevant control component. Thus, the global configuration of $\mathcal{M}$ is not stored in a single state but distributed across all reachable states of $\mathcal{A}$.
Simulation principle.
During the $\omega$-explorability game, Spoiler chooses the next input letter $(t,i)$, specifying which transition of $\mathcal{M}$ to apply and where, while Determiniser responds by moving tokens in $\mathcal{A}$. The automaton $\mathcal{A}$ is constructed so that Spoiler selects transitions of $\mathcal{M}$ when in existential control states, through his choice of letters, whereas Determiniser selects transitions in universal states, by moving to a state $A_{t}$ of the automaton forcing Spoiler to play a transition $t$ next. The component of $\mathcal{A}$ that updates the configuration of $\mathcal{M}$ according to these choices is deterministic.
Iterating runs.
So far, we have described how a single run of $\mathcal{M}$ is simulated. In the reduction, we actually iterate such runs: every time an accepting configuration is reached, Spoiler may play a special letter that sends some tokens to a designated “target” state $\bot$, and restarts the simulation with the remaining tokens. In [BDG+19], this mechanism allowed Spoiler to eliminate tokens one by one, ensuring a win in the population control problem (with finitely many tokens) if and only if he had a winning strategy as $\exists$ in the ATM game.
Novelty: targeting specific tokens.
In our setting, merely accumulating tokens in $\bot$ is insufficient: to win the $\omega$-explorability game, Spoiler must be able to send any specific token to $\bot$ whenever he reaches an accepting configuration of $\mathcal{M}$. We therefore modify the construction of [BDG+19] to allow such targeted elimination, taking into account the different components of the automaton $\mathcal{A}$ that may contain the token to eliminate. Thanks to this refinement, if $\exists$ has a winning strategy in the ATM game, then Spoiler can use it to repeatedly target and eliminate any token, ensuring victory in the $\omega$-explorability game. Conversely, if $\forall$ has a winning strategy, then Determiniser can always either preserve a “leader” token from reaching $\bot$, or reach an accepting sink $\top$ with some token. Thus Determiniser win the $\omega$-explorability game.
Overall, $\mathcal{A}$ is $\omega$-explorable if and only if $\mathcal{M}$ rejects its input, yielding the desired reduction and establishing ExpTime-hardness.
@@ 5.3.2. Detailed proof of Theorem 28
We take an alternating Turing machine
$\mathcal{M}=(\Sigma_{\mathcal{M}},Q_{\mathcal{M}},\Delta_{\mathcal{M}},q_{0}^{\mathcal{M}},q_{f}^{\mathcal{M}})$
with $Q_{\mathcal{M}}=Q_{\exists}\uplus Q_{\forall}$ and $\Delta_{\mathcal{M}}\subseteq Q_{\mathcal{M}}\times\Sigma_{\mathcal{M}}\times Q_{\mathcal{M}}\times\Sigma_{\mathcal{M}}\times\{L,R\}$. It can be seen as a game between two players: existential ($\exists$) and universal ($\forall$). On a given input $w\in(\Sigma_{\mathcal{M}})^{*}$, a play starts in $q_{0}^{\mathcal{M}}$, and the owner of the current state resolves non-determinism by picking a transition compatible with the currently read symbol. The input is accepted iff $\exists$ has a winning strategy. We assume $\mathcal{M}$ uses polynomial space $P(n)$ on inputs of length $n$, i.e., winning strategies avoid configurations with tape longer than $P(n)$.
For simplicity, we assume $\Sigma_{\mathcal{M}}=\{0,1\}$ and that $\mathcal{M}$ alternates between $Q_{\exists}$ and $Q_{\forall}$, starting in $Q_{\exists}$ (so transitions alternate $Q_{\exists}\to Q_{\forall}\to Q_{\exists}\to\cdots$). In our reduction, this means the choice of the next transition is given alternately to Spoiler (for $\exists$) and Determiniser (for $\forall$).
Safety automaton.
We build a safety automaton $\mathcal{A}=(Q,\Sigma,q_{0},\Delta,\bot)$ where a run is accepting iff it never reaches the rejecting sink $\bot$. Let
$Q\;=\;Q_{\mathcal{M}}\;\uplus\;\textit{Pos}\;\uplus\;\textit{Mem}\;\uplus\;\textit{Trans}\;\uplus\;\{q_{0},\texttt{store},\bot,\top\},$
with the components:
Pos $\;=\;[1,P(n)]$ (current head position)
Mem $\;=\;\{\,m_{b,i}\mid b\in\{0,1\},\,i\in[1,P(n)]\,\}$ (tape content at cell $i$ is $b$)
Trans $\;=\;\{E\}\;\cup\;\{A_{t}\mid t\in\Delta_{\mathcal{M}}\}$ (alternation gadget: players’ choices)
and alphabet
$\Sigma\;=$ $\{\,a_{t,p}\mid t\in\Delta_{\mathcal{M}},\;p\in[1,P(n)]\,\}\;\uplus\;\{\texttt{init},\texttt{end}\}$
$\;\uplus\;\{\texttt{check}_{q}\mid q\in Q_{\mathcal{M}}\}\;\uplus\;\{\texttt{check}_{b,i}\mid b\in\{0,1\},\,i\in[1,P(n)]\}.$
Here $\bot$ is a rejecting sink state: a run is accepting if and only if it never reaches this state.
@@
We now define the transitions in $\Delta$. The states $\top$ and $\bot$ are both sinks ($\top$ accepting and $\bot$ rejecting). We then describe all transitions labelled by the letter $a_{t,p}$ with $p\in\textit{Pos}$ and $t=(q,b,q^{\prime},b^{\prime},d)\in\Delta_{\mathcal{M}}$, where $q$ and $q^{\prime}$ are the starting and destination states of $t$, $b$ and $b^{\prime}$ are the letters read and written at the current head position, and $d\in\{L,R\}$ is the direction taken by the head. We give below the list of transitions induced by a letter $a_{t,p}$, grouped by purpose:
Update the configuration of $\mathcal{M}$:
  • $q\rightarrow q^{\prime}$.  (update control state)
  • $p\rightarrow p^{\prime}$ with $p^{\prime}=p+1$ if $d=R$, or p1p-1 if $d=L$. Go to $\top$ if $p^{\prime}\notin[1,P(n)]$.  (update head position)
  • $m_{b,p}\rightarrow m_{b^{\prime},p}$, and $m_{b^{\prime\prime},p^{\prime\prime}}\rightarrow m_{b^{\prime\prime},p^{\prime\prime}}$ for any $b^{\prime\prime}$ and any $p^{\prime\prime}\neq p$.  (update tape content)
Manage alternation between players:
  • $E\rightarrow A_{t^{\prime}}$ for any transition $t^{\prime}$.  (switch to Determiniser’ turn)
  • $A_{t}\rightarrow E$.  (switch to Spoiler’s turn)
Punish invalid choices:
  • $q^{\prime\prime}\rightarrow\top$ for any $q^{\prime\prime}\neq q$.  (punish invalid control state)
  • $m_{\neg b,p}\rightarrow\top$, where $\neg b$ is not $b$.  (punish invalid tape content)
  • $p^{\prime}\rightarrow\top$ for any $p^{\prime}\neq p$.  (punish invalid head position)
Store tokens of non-matching transitions:
@@
These transitions are represented in Figure 5.
@@
Figure 5. Transitions for $a_{t,p}$, where $t=(q,b,q^{\prime},b^{\prime},d)$, $p^{\prime}$ is the position at direction $d$ from $p$, and $q^{\prime\prime}$, $p^{\prime\prime}$, and $t^{\prime}$ are different from $q,p,t$ respectively.
Some auxiliary leters will be used to handle other cases, such as the check letters which handle the case where Determiniser is the one choosing an invalid transition, by sending token to a state $A_{t}$ when transition $t$ is not compatible with the current configuration of $\mathcal{M}$ stored in components $Q_{\mathcal{M}}$, Pos and Mem. We list below all transitions induce by other letters of the alphabet (those not of the form $a_{t,p}$):
  • Initialise configuration of $\mathcal{M}$: init labels transitions from $q_{0}$ to the states $E$, $q_{0}^{\mathcal{M}}$, and $1\in\textit{Pos}$, and also to the states $m_{b,p}$ corresponding to the initial content of the tape, i.e. all $m_{b,p}$ such that $b$ is the $p$-th letter of $w$ (or 0 if $i>|w|$).
  • End current run: end labels transitions from any non-accepting state of $\mathcal{M}$ to $\top$, from store to $q_{0}$, and from any other state to $\bot$.
  • Punish inconsistent choice of $A_{t}$ from Determiniser:
    • $\texttt{check}_{q}$ labels a transition from $A_{t}$ to $\bot$ for any $t\in\Delta$ starting from $q$. It also labels a transition from $q$ to $\top$. Any other state is sent back to $q_{0}$ upon reading $\texttt{check}_{q}$. Intuitively, playing that letter means that $q$ is not the current state and that any transition starting from $q$ is invalid.
    • $\texttt{check}_{b,p}$ labels a transition from $A_{t}$ to $\bot$ for any $t\in\Delta$ reading $b$ on the tape. It also labels transitions from any $j\in\textit{Pos}\setminus\{p\}$ and from $m_{b,p}$ to $\top$. Any other state is sent to $q_{0}$ $\texttt{check}_{b,p}$. Intuitively, playing that letter means that the current head position is $p$, and that its content is not $b$, so any transition reading $b$ is invalid.
To summarize, the states of $\mathcal{A}$ can be seen as two blocks, apart from $q_{0}$, $\top$ and $\bot$: those dealing with the configuration of $\mathcal{M}$ ($Q_{\mathcal{M}}$, Pos and Mem), and those from the gadget of Figure 6, dealing with the alternation and non-deterministic choices.
@@
Expected play pattern.
Let us first describe what a normal play looks like, assuming both players play according to consistent strategies $\sigma_{\exists}$ and $\sigma_{\forall}$ in the game of $\mathcal{M}$ on $w$. This gives a simplified view of the expected behaviour of the players in the $\omega$-explorability game on $\mathcal{A}$.
  • The play starts by Spoiler playing init, which places tokens in the states representing the initial configuration of $\mathcal{M}$ on $w$, along with a target token (the one with minimal available index) in $E$.
  • From there, the players alternate choosing transitions of $\mathcal{M}$:
    • when the target token is in $E$, Spoiler chooses a transition $t$ according to $\sigma_{\exists}$ and plays $a_{t,p}$ where $p$ is the current head position;
    • then Determiniser moves the token to $A_{t^{\prime}}$ where $t^{\prime}=\sigma_{\forall}(\ldots)$, and Spoiler acknowledges his choice by playing $a_{t^{\prime},p^{\prime}}$, where $p^{\prime}$ is the updated head position, and so on from $E$ again. Other tokens than the main one may be located in other states $A_{t^{\prime\prime},p^{\prime}}$, corresponding to other valid transitions $t^{\prime\prime}$ from the target configuration. These ones will be sent to store when $a_{t^{\prime},p^{\prime}}$ is played.
  • When an accepting configuration of $\mathcal{M}$ is reached, Spoiler plays end, which sends all non-store tokens (in particulare the target one) to $\bot$. Tokens in store are sent back to $q_{0}$, and the play can restart with another init.
@@
Lemmas 30 and 31 below ensure the correctness of the reduction.
Lemma 30.
If $w\in\mathcal{L}(\mathcal{M})$, then Spoiler has a winning strategy in the $\omega$-explorability game on $\mathcal{A}$.
Proof 5.6.
There is a winning strategy $\sigma_{\exists}$ for $\exists$ in the alternating Turing machine game, and Spoiler will use that strategy in the explorability game to win against $\omega$ tokens. He will consider that the tokens are labelled by integers, and will always target the smallest one that is not already in $\bot$. He proceeds as follows.
@@
  • After such a play, Determiniser can move tokens to any state $A_{t}$. If he chooses several such states, then Spoiler will only pay attention to the one containing the current target token.
    • If that state $A_{t}$ corresponds to an invalid transition $t$ (wrong starting state or wrong tape content at the current head position), then Spoiler plays the corresponding check letter. Formally, if the target token is in $\mathcal{A}_{t}$, Spoiler plays $\texttt{check}_{q}$ if the starting state $q$ of $t$ does not match the current state of the tape (given by Lemma 29), or $\texttt{check}_{b,p}$ if the current head position is $p$ and does not contain $b$. In both cases, the target token is sent to $\bot$ with no other token reaching $\top$ (by Lemma 29). This sends us back to the first step, but with an updated target.
    • If $t$ is consistent with the current configuration, then Spoiler can play the corresponding $a_{t,p}$, where $p$ is the current head position (again, given by Lemma 29), then go back to the previous step (where there are tokens in $E$).
  • If the run proceeds normally with consistent transitions, it will eventually reach an accepting state of $\mathcal{M}$ because $\sigma_{\exists}$ is winning. This corresponds to a stage where Spoiler can safely play end to send the target token to $\bot$ along with all tokens outside of store. This sends us back to the first step, but with an updated target. Notice that if there was a virtual target token, we will always reach this event, and send the real target token (located in $Q_{\mathcal{M}}$ or Pos or Mem) in $\bot$.
@@
Lemma 31.
If $w\notin\mathcal{L}(\mathcal{M})$, then Determiniser has a winning strategy in the $\omega$-explorability game on $\mathcal{A}$.
Proof 5.7.
We now assume that there is a winning strategy $\sigma_{\forall}$ for the universal player in the alternation game on $\mathcal{M}(w)$, and we build a winning strategy for Determiniser in the $\omega$-explorability game. This strategy is more straightforward than the previous one, as we can focus on the tokens sent to $E$ (while still populating each state when init is played, but these other tokens follow a deterministic path until the next init).
Determiniser will initially choose a specific token, called leader. He then sends $\omega$ tokens to every reachable state when Spoiler plays init, with the leader going to $E$. Determiniser then moves leader according to $\sigma_{\forall}$, and always keeps every reachable state populated by $\omega$ tokens. Spoiler cannot send the leader to $\bot$, since the only way to do that would be using the letter end, but this would immediately ensure the win for Determiniser, as there will always be some token in non-accepting states of $\mathcal{M}$ (because $\sigma_{\forall}$ is winning), and those tokens would be sent to $\top$ upon playing end. This means that Spoiler has no way to send the leader to $\bot$ without losing the game, and therefore Determiniser wins.
Note that with that strategy, Spoiler can still send some tokens to $\bot$ during each run. However, Determiniser will start the next run with still $\omega$ tokens, including the leader. This is why the choice of a specific leader is important, as this is the witness that an accepting run will be built by $Determiniser$.